OpenStudy (anonymous):

Use the definition of the derivative (don't be tempted to take shortcuts!) to find the derivative of the function f(x)=7−8x+7x^2. Then state the domain of the function and the domain of the derivative. Note: When entering interval notation in WeBWorK, use I for ∞, -I for −∞, and U for the union symbol. If the set is empty, enter "{}" without the quotation marks.

OpenStudy (anonymous):

got f'(x) but cant get domains

OpenStudy (anonymous):

@Algebraic! could you help me?

OpenStudy (anonymous):

14x-8 is derivative

OpenStudy (anonymous):

\[\lim_{h \rightarrow0} \frac{ f(x+h) -f(x) }{h }\]

OpenStudy (anonymous):

\[\lim_{h \rightarrow0} \frac{ 7(x+h)^2 -8(x+h) +7 -(7x^2 -8x +7) }{h }\]

OpenStudy (anonymous):

expand (x+h)^2 and then simplify the expression...

OpenStudy (anonymous):

when it's simplified so that there's no longer a problem with taking the limit as h->0, then take the limit by substitution.

OpenStudy (anonymous):

is this the way you did it?

OpenStudy (anonymous):

ya and i got x=4/7

OpenStudy (anonymous):

14x-8 ...?

OpenStudy (anonymous):

try it; you should get the derivative.. since that's what this expression is...

OpenStudy (anonymous):

ya i got 14x-8 for the derivitave

OpenStudy (anonymous):

ok, the domain is all x

OpenStudy (anonymous):

ya \[x \neq4/7\]

OpenStudy (anonymous):

why's that?

OpenStudy (anonymous):

(-inf,4/7)U(4/7,inf)

OpenStudy (anonymous):

because if you plug that in y=0

OpenStudy (anonymous):

that would make it a root. it's not a root of either the function or the derivative though, so not sure where you're getting it. roots are part of the domain anyway... I really don't have any idea what you did so far on this problem... did you use the limit definition to find the derivative?

OpenStudy (anonymous):

idk im super confused

OpenStudy (anonymous):

the function is a parabola, so there's no restriction on x values it can have... the derivative is a line so there's no restriction on x values it can have either...

OpenStudy (anonymous):

ok so how i ind domain of fx